# A Survey on Large Language Model Acceleration based on KV Cache Management

### v2

@@ 6.2.2 Preemptive and Fairness-oriented scheduling
FastSwitch [225] introduces a fairness-oriented KV cache scheduling system that addresses the overhead challenges of preemptive scheduling in LLM serving. There are three key mechanisms: enhancing I/O utilization through intelligent cache movement scheduling, minimizing GPU idle time during context switches, and eliminating redundant I/O operations in multi-turn conversations. Unlike traditional block-based KV cache memory policies that prioritize memory efficiency at the cost of fragmentation and granularity limitations, FastSwitch [225] implements a balanced approach that maintains efficient memory usage while facilitating smoother context switching. This integrated scheduling approach enables dynamic priority adjustments for fairness while minimizing the performance impact of context switches.
@@ 6.2.3 Layer-specific and Hierarchical Scheduling
LayerKV [232] introduces a novel layer-wise KV cache scheduling approach to address the growing TTFT (Time to First Token) latency challenges in large-context LLM serving. The contribution lies in its fine-grained, layer-specific KV cache block allocation and management strategy, which departs from traditional monolithic cache management approaches. By implementing layer-wise KV block scheduling and offloading mechanisms, LayerKV [232] enables more efficient memory utilization and reduces queuing delays that typically occur when large context windows compete for limited GPU KV cache blocks. It is complemented by an SLO-aware scheduler that optimizes cache allocation decisions based on service level objectives, allowing for dynamic management of memory resources across model layers.
@@
ALISA [234] introduces a dual-level KV cache scheduling framework that combines algorithmic sparsity with system-level optimization. At the algorithm level, the Sparse Window Attention mechanism identifies and prioritizes the most important tokens for attention computation, creating a mixture of global dynamic and local static sparse patterns that significantly reduces KV cache memory requirements. At the system-level, its three-phase token-level dynamic scheduler that manages KV tensor allocation and optimizes the trade-off between caching and recomputation. The scheduler makes dynamic decisions about which tokens to cache in GPU memory versus recompute, based on their importance and system resource constraints.
@@ 6.2.4 Summary and Future Directions
Future work should advance these KV cache scheduling innovations through several interlinked dimensions: developing adaptive hybrid systems that dynamically select optimal scheduling strategies based on real-time workload characteristics, exploring predictive models that anticipate user request patterns to proactively optimize cache allocation, investigating automated parameter tuning mechanisms that adjust scheduling policies across different deployment scenarios, designing context-aware architectures that intelligently balance prefix sharing with fairness requirements, and researching novel cache coherence protocols that efficiently handle distributed inference scenarios. Additionally, the integration of reinforcement learning approaches could enable self-optimizing schedulers that learn from historical usage patterns, while the exploration of hardware-software co-design could yield specialized accelerators that directly support efficient KV cache management operations. These advancements would collectively enhance the robustness, efficiency, and adaptability of LLM inference systems across diverse operational conditions and deployment scales. Finally, considering LLM serving [240], different scheduling and sharing for multiple users and queries may lead to potential privacy leaks. Therefore, privacy protection techniques for LLM serving in multi-user scenarios, such as differential privacy [241, 242, 243], are worth further investigation.
@@
Method Single/Multi-GPU I/O-aware Heterogeneous SSD-based
@@ 6.3.1 Single/Multi-GPU Design
Based on these works focusing on GPU-oriented designs, we can categorize the approaches into several key strategies for KV cache optimization. First, shared prefix optimization approaches like HydraGen [226] and DeFT [227] focus on efficient GPU memory utilization through batched prefix computations and tree-structured attention patterns. Rather than maintaining separate KV caches for each sequence with identical prefixes, HydraGen [226] decomposes attention computation to leverage a single shared KV cache for common prefixes across multiple requests. It enables efficient GPU memory utilization through two mechanisms: batched prefix KV cache access across sequences and separate handling of unique suffix KV caches. For DeFT [227], its core contributions are twofold: KV-Guided Grouping, which optimizes GPU memory access patterns by intelligently managing shared prefix KV caches to minimize redundant global-to-shared memory transfers, and Flattened Tree KV Splitting, which ensures balanced workload distribution across GPU compute units while minimizing computational redundancy.
@@
Third, phase-aware designs like DistServe [229] separate prefill and decoding phases across GPU resources to optimize their distinct memory access patterns. Novel batching strategies are represented by Multi-Bin Batching [230], which focuses on length-aware request grouping for improved GPU utilization, while advanced parallel computation frameworks like Tree Attention [231] introduce sophisticated reduction algorithms for efficient attention computation across multiple GPUs. DistServe [229] recognizes that prefill and decoding phases have distinct KV cache utilization characteristics and memory access patterns: prefill requires intensive computation with growing KV cache sizes for processing input tokens, while decoding maintains a fixed KV cache size for generating output tokens. By physically separating these phases onto different GPUs, DistServe enables optimized GPU memory management and KV cache access patterns specific to each phase, eliminating interference between prefill’s bursty memory access patterns and decoding’s steady-state KV cache utilization. Multi-Bin Batching [230] introduces a length-aware batching strategy helps minimize GPU idle time and memory fragmentation that typically occurs when processing requests of varying lengths in the same batch, as it ensures that the KV cache memory allocated for each batch is utilized more uniformly across all requests. Tree Attention [231] implements a tree-based reduction algorithm that fundamentally changes how attention values are computed and aggregated across GPUs, enabling more efficient handling of KV cache data through partial reductions that significantly reduce memory bandwidth requirements and peak memory usage.
@@ 6.3.2 I/O-based Design
Recent I/O-focused optimizations for KV cache management span several key dimensions, targeting different levels of the memory hierarchy. At the GPU level, approaches like FlashAttention [145] and Bifurcated Attention [222] optimize data movement between HBM and SRAM through sophisticated tiling strategies and split attention computations, while CPU-GPU data movement optimizations are addressed by systems like PartKVRec [221], which tackles PCIe bandwidth bottlenecks through hybrid recomputation and transfer strategies, and HCache [223], which optimizes intermediate activation storage and restoration.
FlashAttention [145] employs a tiling strategy that carefully manages KV cache access patterns, reducing redundant memory operations by keeping frequently accessed portions of the KV cache in fast SRAM while systematically fetching and evicting data blocks to minimize HBM accesses. Bifurcated Attention [222] presents an I/O-aware approach to optimize KV cache access patterns during shared-context batch decoding by strategically splitting attention computations into two distinct GEMM operations. It specifically targets the memory bandwidth bottleneck in high-batch scenarios with long contexts by minimizing repeated KV cache accesses, maintaining the same computational FLOPs while drastically reducing memory I/O operations. For PartKVRec [221], its key innovation lies in its hybrid strategy of partial KV cache recomputation on the GPU while simultaneously transferring the remaining cache data from CPU memory, effectively hiding PCIe transfer latency. The implementation employs a sophisticated I/O-aware scheduling system that analyzes input characteristics and hardware capabilities to determine the optimal balance between recomputation and data transfer, dynamically managing KV cache movement to maximize PCIe bandwidth utilization while minimizing GPU idle time. HCache [223] strategically stores and restores intermediate activations instead of complete KV cache states, implementing a bubble-free restoration scheduler that carefully balances computation and I/O operations to maximize bandwidth utilization. A key innovation is its chunk-based storage manager that addresses the I/O pattern mismatch between saving (layer-before-token) and restoration (token-before-layer) operations, optimizing data layout and access patterns to reduce I/O overhead. Cake [224] addresses the fundamental I/O bottleneck in loading cached KV states from disk to GPU memory. It introduces a bidirectional parallelized strategy that simultaneously leverages both computational and I/O resources. This hybrid approach dynamically balances between loading cached KV states from storage and computing them on GPUs, adapting automatically to varying system conditions without manual parameter tuning.
@@ 6.3.3 Heterogeneous Design
Recent heterogeneous computing approaches for KV Cache demonstrate diverse strategies for optimizing CPU-GPU collaboration. Systems like NEO [216] and FastDecode [217] implement strategic workload distribution through CPU offloading of attention computations, while FlexInfer [218] introduces virtual memory abstractions for optimal resource coordination.
NEO [216] advances heterogeneous computing for LLM inference by implementing strategic CPU offloading of attention computations and KV cache states. Through asymmetric GPU-CPU pipelining and load-aware scheduling, it optimally balances workloads across both computing platforms, enabling larger GPU batch sizes without latency penalties. For FastDecode [217], its key contribution lies in its strategic offloading of memory-bound KV cache operations to distributed CPU resources, leveraging the aggregate memory capacity and computing power of multiple CPU nodes rather than treating CPUs as mere storage devices. By utilizing CPUs for KV cache computations and storage while keeping compute-intensive operations on GPUs, it creates an efficient pipeline that maximizes resource utilization across the heterogeneous infrastructure, enabling larger batch sizes and higher throughput. FlexInfer [218] orchestrates CPU-GPU resource utilization for LLM inference by introducing the virtual memory-based abstraction vTensor.
@@ 6.3.5 Summary and Future Directions
Recent advancements in hardware-aware designs for KV cache management emphasize optimizing performance based on specific hardware architectures and constraints, demonstrating significant enhancements in large language model inference efficiency. Approaches like HydraGen [226] and vLLM [144] in single and multi-GPU designs focus on efficient memory access patterns and load balancing, while I/O-based strategies such as FlashAttention [145] and PartKVRec [221] tackle data movement bottlenecks through intelligent prefetching and scheduling mechanisms. Additionally, heterogeneous designs exemplified by NEO [216] and FastDecode [217] effectively leverage CPU-GPU collaboration to maximize resource utilization.
Future work should advance this research through multiple interconnected directions: exploring novel architectural designs that combine specialized hardware accelerators with optimized memory hierarchies, investigating hybrid systems that leverage computational storage drives and processing-in-memory capabilities, developing self-adaptive algorithms that dynamically optimize resource allocation based on workload patterns, researching advanced compression techniques that maintain model fidelity while reducing memory requirements, and designing intelligent scheduling mechanisms that efficiently coordinate heterogeneous computing resources including CPUs, GPUs, and custom accelerators. These improvements, working in concert, would enhance both the performance and scalability of LLM inference systems across diverse deployment scenarios, from edge devices to data centers, while maintaining adaptability to emerging hardware innovations and varying computational demands.
7 Text and Multi-modal Datasets
@@
7.1 Text Dataset
We collect a lot of long-context datasets from state-of-the-art benchmark frameworks and various papers, including L-Eval[244], M4LE[245], BAMBOO[246], LongBench[247], LRA[248], SCROLLS[249], ZEROSCROLLS[250], LooGLE[251], LongEval[252], and StreamingEval[136]. Specifically, we categorize these datasets into different tasks, including question answering, text summarization, text reasoning, text retrieval, and text generation.
TABLE XIII: Text question answering (QA) dataset. In the Avg. Len: average length, Tok: tokens; W: words. In the Instances column, Doc: documents, Q: questions, Inst: instructions. Particularly, AltQA, PaperQA and MeetingQA have two datasets with different length levels, and is separated with ///.
Particularly, for datasets from L-Eval, the GPT-4 metric means the win-rate against Turbo-16K, judged by GPT-4. $\Delta L$ is the length difference between answer length and ground truth. For NarrativeQA, MRR: Mean Reciprocal Rank .
Task Name Source Instances Avg Len Metric Lang.
QA AltQA [253] Wikipedia 200/200 3243/13,084 Tok Acc EN
QA PaperQA(BAMBOO) [246] Paper 100/100 3101/6838 Tok Acc EN
QA MeetingQA(BAMBOO [246] Meeting 100/100 2738/9838 Tok Acc EN
QA TriviaQA [254] Web Question, Wiki 95,956 Q, 662,659 Doc 17,370 W EM, F1 EN
QA TOEFL(L-Eval) [244] TOFEL-QA[255] 15 Doc, 269 Inst 3907 Tok Rouge-L, GPT-4, $\Delta L$ EN
QA Coursera(L-Eval) [244] Video Subtitles 15 Doc, 172 Inst 9075 Tok Rouge-L, GPT-4, $\Delta L$ EN
QA SFiction(L-Eval) [244] SFGram [256], fiction 7 Doc, 64 Inst 16,381 Tok Rouge-L, GPT-4, $\Delta L$ EN
QA LongFQA(L-Eval) [244] Financial Transcripts 6 Doc, 52 Inst 6032 Tok Rouge-L, GPT-4, $\Delta L$ EN
QA CUAD(L-Eval) [244] CUAD[257] 20 Doc, 130 Inst 30,966 Tok Rouge-L, GPT-4, $\Delta L$ EN
QA DuoRC [245] Movie 3572 W Acc EN
QA NQ [258] Wiki 307,373 9005 W Rouge EN
QA-SG NarrativeQA [259] Story 1572 Doc 62,528 Tok
BLEU, METEOR,
Rouge-L, MRR

### v3

@@ 6.2.2 Preemptive and Fairness-oriented scheduling
FastSwitch [238] innovatively introduces a fairness-oriented KV cache scheduling system that addresses the overhead challenges of preemptive scheduling in LLM serving. There are three key mechanisms: enhancing I/O utilization through intelligent cache movement scheduling, minimizing GPU idle time during context switches, and eliminating redundant I/O operations in multi-turn conversations. Unlike traditional block-based KV cache memory policies that prioritize memory efficiency at the cost of fragmentation and granularity limitations, FastSwitch [238] implements a balanced approach that maintains efficient memory usage while facilitating smoother context switching. This integrated scheduling approach enables dynamic priority adjustments for fairness while minimizing the performance impact of context switches.
FlowKV [254] introduces the Load-Aware Scheduler for balanced request scheduling and flexible PD node allocation. This design maximizes hardware resource utilization, achieving peak system throughput across various scenarios, including normal, computational imbalance, and extreme overload conditions.
@@ 6.2.3 Layer-specific and Hierarchical Scheduling
LayerKV [248] innovatively introduces a novel layer-wise KV cache scheduling approach to address the growing TTFT (Time to First Token) latency challenges in large-context LLM serving. The contribution lies in its fine-grained, layer-specific KV cache block allocation and management strategy, which departs from traditional monolithic cache management approaches. By implementing layer-wise KV block scheduling and offloading mechanisms, LayerKV [248] enables more efficient memory utilization and reduces queuing delays that typically occur when large context windows compete for limited GPU KV cache blocks. It is complemented by an SLO-aware scheduler that optimizes cache allocation decisions based on service level objectives, allowing for dynamic management of critical memory resources across model layers.
@@
ALISA [250] introduces a dual-level KV cache scheduling framework that combines algorithmic sparsity with system-level optimization. At the algorithm level, the Sparse Window Attention mechanism identifies and prioritizes the most important tokens for attention computation, creating a mixture of global dynamic and local static sparse patterns that significantly reduce KV cache memory requirements. At the system-level, its three-phase token-level dynamic scheduler that manages KV tensor allocation and optimizes the trade-off between caching and recomputation. The scheduler makes dynamic decisions about which tokens to cache in GPU memory versus recompute, based on their importance and system resource constraints.
@@
Apt-Serve [252] employs an adaptive runtime scheduling mechanism that dynamically optimizes batch composition based the hybrid cache. It defines the adaptive scheduling optimization problem and proposes an efficient algorithm with theoretical guarantees.
FGOS [253] presents an approach to optimizing large language model (LLM) inference by formulating it as a multi-stage online scheduling problem. It recognize that the sequential arrival of prompts and the growth of the KV cache render conventional scheduling techniques ineffective in this context. To tackle this challenge, it develops a fluid dynamics approximation that provides a tractable benchmark to guide their algorithm design. Building on this foundation, it utilizes carefully-designed multiple thresholds to schedule incoming prompts optimally when the output lengths are known.
@@ 6.2.4 Summary and Future Directions
Future KV cache scheduling work should focus on adaptive and predictive systems, automated tuning, context-aware architectures, and novel coherence protocols. Integrating reinforcement learning and hardware-software co-design will further enhance LLM inference system robustness, efficiency, and adaptability. Finally, considering LLM serving [263], different scheduling and sharing for multiple users and queries may lead to potential privacy leaks. Therefore, privacy protection techniques for LLM serving in multi-user scenarios, such as differential privacy [264, 265, 266], are worth further investigation.
@@
Method
Single/Multi-GPU
I/O-aware
Heterogeneous
SSD-based
APEX [234]
@@ 6.3.1 Single/Multi-GPU Design
Based on these comprehensive works focusing on GPU-oriented designs, we can categorize the approaches into several key strategies for KV cache optimization. First, shared prefix optimization approaches like HydraGen [239] and DeFT [240] focus on efficient GPU memory utilization through batched prefix computations and tree-structured attention patterns. Rather than maintaining separate KV caches for each sequence with identical prefixes, HydraGen [239] decomposes attention computation to leverage a single shared KV cache for common prefixes across multiple requests. It enables efficient GPU memory utilization through two mechanisms: batched prefix KV cache access across sequences and separate handling of unique suffix KV caches. For DeFT [240], its core contributions are twofold: KV-Guided Grouping, which optimizes GPU memory access patterns by intelligently managing shared prefix KV caches to minimize redundant global-to-shared memory transfers, and Flattened Tree KV Splitting, which ensures balanced workload distribution across GPU compute units while minimizing computational redundancy.
@@
Third, phase-aware designs like DistServe [242] separate prefill and decoding phases across GPU resources to optimize their distinct memory access patterns. Novel batching strategies are represented by Multi-Bin Batching [243], which focuses on length-aware request grouping for improved GPU utilization, while advanced parallel computation frameworks like Tree Attention [244] introduce sophisticated reduction algorithms for efficient attention computation across multiple GPUs. DistServe [242] recognizes that prefill and decoding phases have distinct KV cache utilization characteristics and memory access patterns: prefill requires intensive computation with growing KV cache sizes for processing input tokens, while decoding maintains a fixed KV cache size for generating output tokens. By physically separating these phases onto different GPUs, DistServe enables optimized GPU memory management and KV cache access patterns specific to each phase, eliminating interference between prefill’s bursty memory access patterns and decoding’s steady-state KV cache utilization. Multi-Bin Batching [243] introduces a length-aware batching strategy that helps minimize GPU idle time and memory fragmentation that typically occurs when processing requests of varying lengths in the same batch, as it ensures that the KV cache memory allocated for each batch is utilized more uniformly across all requests. Tree Attention [244] implements a tree-based reduction algorithm that fundamentally changes how attention values are computed and aggregated across GPUs, enabling more efficient handling of KV cache data through partial reductions that significantly reduce memory bandwidth requirements and peak memory usage.
Recently, gLLM [245] is a globally balanced pipeline parallelism system incorporating Token Throttling to effectively mitigate the pipeline bubbles. Its runtime adopts an asynchronous execution and message-passing architecture specifically optimized for pipeline parallelism characteristics. FairKV [246] is a method designed to ensure fair memory usage among attention heads in systems employing imbalanced KV cache compression. The core technique of FairKV is Fair-Copying, which replicates a small subset of memory-intensive attention heads across GPUs using data parallelism to mitigate load imbalance. MELL [247] is a memory-efficient LLM serving system via multi-GPU KV cache management. It saves the number of GPUs needed in the system by considering the dynamic KV cache load and the costly request migration.
@@ 6.3.2 I/O-based Design
Recent I/O-focused optimizations for KV cache management span several key dimensions, targeting different levels of the memory hierarchy. At the GPU level, approaches like FlashAttention [150] [151] [152] and Bifurcated Attention [235] optimize data movement between HBM and SRAM through sophisticated tiling strategies and split attention computations, while CPU-GPU data movement optimizations are addressed by systems like PartKVRec [232], which tackles PCIe bandwidth bottlenecks through hybrid recomputation and transfer strategies, and HCache [236], which optimizes intermediate activation storage and restoration.
FlashAttention [150] [151] [152] employs a tiling strategy that carefully manages KV cache access patterns, reducing redundant memory operations by keeping frequently accessed portions of the KV cache in fast SRAM while systematically fetching and evicting data blocks to minimize HBM accesses. Bifurcated Attention [235] presents an I/O-aware approach to optimize KV cache access patterns during shared-context batch decoding by strategically splitting attention computations into two distinct GEMM operations. It specifically targets the memory bandwidth bottleneck in high-batch scenarios with long contexts by minimizing repeated KV cache accesses, maintaining the same computational FLOPs while drastically reducing memory I/O operations. For PartKVRec [232], its key innovation lies in its hybrid strategy of partial KV cache recomputation on the GPU while simultaneously transferring the remaining cache data from CPU memory, effectively hiding PCIe transfer latency. The implementation employs a sophisticated I/O-aware scheduling system that analyzes input characteristics and hardware capabilities to determine the optimal balance between recomputation and data transfer, dynamically managing KV cache movement to maximize PCIe bandwidth utilization while minimizing GPU idle time. HCache [236] strategically stores and restores intermediate activations instead of complete KV cache states, implementing a bubble-free restoration scheduler that carefully balances computation and I/O operations to maximize bandwidth utilization. A key innovation is its chunk-based storage manager that addresses the I/O pattern mismatch between saving (layer-before-token) and restoration (token-before-layer) operations, optimizing data layout and access patterns to reduce I/O overhead. Cake [237] addresses the fundamental I/O bottleneck in loading cached KV states from disk to GPU memory. It introduces a bidirectional parallelized strategy that simultaneously leverages both computational and I/O resources. This hybrid approach dynamically balances between loading cached KV states from storage and computing them on GPUs, adapting automatically to varying system conditions without manual parameter tuning.
@@ 6.3.3 Heterogeneous Design
Recent heterogeneous computing approaches for KV Cache demonstrate diverse strategies for optimizing CPU-GPU collaboration. Systems like NEO [227], FastDecode [228] and HeadInfer [233] implement strategic workload distribution through CPU offloading of attention computations, while FlexInfer [229] introduces virtual memory abstractions for optimal resource coordination.
NEO [227] advances heterogeneous computing for LLM inference by implementing strategic CPU offloading of attention computations and KV cache states. Through asymmetric GPU-CPU pipelining and load-aware scheduling, it optimally balances workloads across both computing platforms, enabling larger GPU batch sizes without latency penalties. For FastDecode [228], its key contribution lies in its strategic offloading of memory-bound KV cache operations to distributed CPU resources, leveraging the aggregate memory capacity and computing power of multiple CPU nodes rather than treating CPUs as mere storage devices. By utilizing CPUs for KV cache computations and storage while keeping compute-intensive operations on GPUs, it creates an efficient pipeline that maximizes resource utilization across the heterogeneous infrastructure, enabling larger batch sizes and higher throughput. FlexInfer [229] orchestrates CPU-GPU resource utilization for LLM inference by introducing the virtual memory-based abstraction vTensor. By implementing fine-grained, selective offloading of attention heads’ KV cache to CPU RAM while dynamically computing attention outputs, HeadInfer [233] achieves remarkable memory efficiency without compromising computational performance. It also supports both dense and sparse attention mechanisms, integrates with pipeline parallelism for larger models. Unlike NEO [227], which focuses on strategic CPU offloading with asymmetric GPU-CPU pipelining, and FastDecode [228], which distributes KV cache operations across multiple CPU nodes, HeadInfer distinguishes itself through its granular head-wise offloading strategy that eliminates the need to fully store KV cache for any transformer layer on GPU while preserving complete mathematical equivalence without approximation.
@@
APEX [234] is a profiling-informed scheduling strategy that maximizes CPU-GPU parallelism during hybrid LLM inference. Unlike systems relying on static rules or purely heuristic approaches, APEX dynamically dispatches compute across heterogeneous resources by predicting execution times of CPU and GPU subtasks to maximize overlap while avoiding scheduling overheads.
@@ 6.3.5 Summary and Future Directions
Recent advancements in hardware-aware designs for KV cache management emphasize optimizing performance based on specific hardware architectures and constraints, demonstrating significant enhancements in large language model inference efficiency. Approaches like HydraGen [239] and vLLM [149] in single and multi-GPU designs focus on efficient memory access patterns and load balancing, while I/O-based strategies such as FlashAttention [150] [151] [152] and PartKVRec [232] tackle data movement bottlenecks through intelligent prefetching and scheduling mechanisms. Additionally, heterogeneous designs exemplified by NEO [227] and FastDecode [228] effectively leverage CPU-GPU collaboration to maximize resource utilization.
TABLE XIII: Long-context Text Benchmarks.
Benchmark Tasks Language
Q-A Summarization Reasoning Retrieval Generation Aggregation
MultiTurnBench [1]
EN
NumericBench [267]
EN
RULER [268]
EN
OneRuler [269]
26 languages
L-Eval [270]
EN
M4LE [271]
EN/ZH
BAMBOO [272]
EN
LongBench [273]
EN/ZH
SCROLLS [274]
EN
ZEROSCROLLS [275]
EN
LooGLE [276]
EN
LongEval [277]
EN
StreamingEval [141]
EN
Future research will focus on enhancing LLM inference systems through interconnected advancements in revolutionary architectural design, hybrid systems leveraging computational storage and in-memory processing, adaptive resource allocation algorithms, advanced compression techniques, and intelligent scheduling for heterogeneous computing. These efforts aim to improve performance and scalability across diverse deployments while adapting to new hardware and demands.
7 Long-context Text and Multi-modal Benchmarks
@@
7.1 Text Benchmarks
We collect a lot of long-context datasets, such as NumericBench [267]and LongBench [273]. We categorize these datasets into various tasks, including question answering, text summarization, text reasoning, text retrieval, text generation, and aggregation. The brief descriptions of each benchmark are provided below, and the corresponding statistics are listed in Tab. XIII
TABLE XIV: Multi-modal Benchmark Tasks. Specifically, for task abbreviation, Conv: conversation task; Desc: description task; Reas: reasoning task; Perc: perception task; Pred: prediction task; SUMM: summary task.
Benchmark Tasks Language
Conv Desc Reas Perc Pred Count Retrieval Order SUMM
LLaVA-Bench [278]